OpenStudy (anonymous):

for which value of k will the pair of linear equations kx+3y=k-3 and 12x+ky=k have no solution?

OpenStudy (amistre64):

thats alot of ks ...

OpenStudy (amistre64):

in effect; when 2 linear equations like that have the same left side, but different right sides; they never are equal

OpenStudy (anonymous):

For no Solution these two lines(equations) shud become parallel to each other,multiply their slopes and equate that to 1 m1*m2=1

OpenStudy (anonymous):

And find k

OpenStudy (amistre64):

kx + 3y = 12x + ky kx + 3y -12x - ky --------------- (k-12)x + (3-k)y = 0 umm, if they have the same slope, them m1*m2 will not equate to 1 in general will it?

OpenStudy (anonymous):

oh,i forgot the condition for parallel lines.m1=m2

OpenStudy (anonymous):

right? @amistre64

OpenStudy (anonymous):

em nt ggttng d answer guyzzzzz

OpenStudy (amistre64):

yes :) or to use your vernacular \(m_1*m_2^{-1}=1\) :)

OpenStudy (amistre64):

what have you tried?

OpenStudy (anonymous):

kx + 3y -12x - ky --------------- (k-12)x + (3-k)y = 0

OpenStudy (anonymous):

after dt???????////

OpenStudy (anonymous):

cn u plzzz show it after dt i will telll where i went wrong

OpenStudy (amistre64):

just an idea; but another might to be to determine when they are both equal to each other kx+3y-k+3 = 12x+ky -k kx-k = 12x -3y -3 +ky -k kx-k -ky +k = 12x -3y -3 k(x -1 -y +1) = 12x -3y -3 \[ k=\frac{12x -3y -3}{x-y}\]

OpenStudy (anonymous):

m1=-k/3, m2=-12/k Find thier slopes by differentiating(its easy) -k/3=-12/k k^2=36 k=+ or - 6

OpenStudy (amistre64):

we are allowed to use calculus? hmmm.....

OpenStudy (amistre64):

im thinking matrix linear algebra stuff in the end :)

OpenStudy (anonymous):

thx guyzzzzzzzzzz

OpenStudy (asnaseer):

first re-arrange both equation in slope-intercept form to get:\[y=-\frac{kx}{3}+\frac{k-3}{3}\tag{1}\]and:\[y=-\frac{12x}{k}+1\tag{2}\]now, for no solutions, both lines must have the same slope but different y-intercepts. If they had the same y-intercept then there are an infinite number of solutions. so, with these conditions we get:\[-\frac{k}{3}=-\frac{12}{k}\implies k^2=36\implies k=\pm6\tag{3}\]and:\[\frac{k-3}{3}\ne1\implies k-3\ne3\implies k\ne 6\tag{4}\]now combine the results of (3) and (4) to find the answer.